23. A car moving at a speed of 72km / h on a straight road, brakes by moving with acceleration a = -2m / s² until it stops. a) After how long will the speed of the car be halved? b) After how long will the car stop? C) How far is it from the moment of braking until it stops?


1
Expert's answer
2021-05-31T08:45:14-0400

a)

"v=v_0+at,""t=\dfrac{v-v_0}{a},""t=\dfrac{10\ \dfrac{m}{s}-20\ \dfrac{m}{s}}{-2\ \dfrac{m}{s^2}}=5\ s."

b)

"t=\dfrac{v-v_0}{a},""t=\dfrac{0-20\ \dfrac{m}{s}}{-2\ \dfrac{m}{s^2}}=10\ s."

c)

"v^2=v_0^2+2ad,""d=\dfrac{v^2-v_0^2}{2a},""d=\dfrac{0-(20\ \dfrac{m}{s})^2}{2\cdot(-2\ \dfrac{m}{s^2})}=100\ m."